OpenStudy (ayonnaleflore):

The height of a triangle is 5 m less than its base. The area of the triangle is 42 m2. Find the length of the base. 12 m 7 m 8 m 11 m

OpenStudy (error1603):

Welcome to OpenStudy!

OpenStudy (error1603):

then altitude = b - 5 then area = 1 / 2 * altitude * base => area = 0.5 ( b - 5 ) b = 42 => b^2 - 5b = 84 => b^2 - 5b - 84 = 0 => ( b - 12 ) ( b + 7 ) = 0 => b - 12 = 0 or b + 7 = 0 => b = 12 or b = -7 since length can't be negative b = 12 so base =12 and altitude = 7

OpenStudy (anonymous):

B. 7 m

OpenStudy (error1603):

Yeah dude

OpenStudy (ayonnaleflore):

ok thanks @Error1603 @EpicGammer

OpenStudy (anonymous):

No problem fan me and let me know if you have more questions?

OpenStudy (ayonnaleflore):

ok i will